OpenStudy (barrelracing):

write an equation in slope intercept form of a line with the given slope m and y-intercept b.

pooja195 (pooja195):

pick a number for slope pick a number for y intercept plug it in y=mx+b

OpenStudy (barrelracing):

oh the numbers are m=-5 and b=7/2. so it would be y=-5x+7/2?